Find the equation of the quadratic function f whose graph is shown below. (-5,7) (-6) 5 ?
Find a polynomial f(x) of degree 3 with real coefficients and the following zeros. 4, 3-i f(x) = x 3 ?
Answer #1

Equation of Quadratic function in vertex form: y = a (2-6)2+ k vertex + (hik) flere, (hik) = (-5,7) ; Given a point = (-6,4);q=3-1 501 degree =3 ler f(x) = a (a-p)(1-2) (2-») Given Zeros, pa 4 rasti found = a (x-4)(3-(3-1))(--(3+;)) ca (6-4) ((2-3)1